A balloon is rising vertically upward with a uniform velocity 'u' as shown At a certain instant, a ball No. 1 is dropped trom it, which takes a 4 seconds to reach the ground. The moment this ball strikes the ground, another ball is dropped from the balloon and this 2nd ball takes 5 seconds to reach the ground. Find the velocity of the balloon 'u' and height from which both the balls must have been dropped.

1
Expert's answer
2022-02-22T11:50:03-0500

The initial velocity with respect to ground will be 10+5=15m/safter 2sec its velocity will be v=15−gt=15−10×2=−5m/sWhose value is 5m/sNote- Once the stone is projected from balloon it will not have any impact of the balloon after that moment.